OpenStudy (anonymous):

find the projection of u along v u=<-1,2,0> and v=<2,0,1>

OpenStudy (zpupster):

the formula for projection: \[\frac{ u \cdot v }{ \left| v \right|^2 }(v)\]

OpenStudy (zpupster):

so \[{ u \cdot v }\] = (-1(2))+2(0)+(0)1) = -2 the magnitude of |v| \[\sqrt{2^2 + 0^2+1^2}\]=\[\sqrt{5}\] so \[\frac{ -2 }{ \sqrt{5} }<2,0,1>\]